# ChartAssisstant: A Universal Chart Multimodal Language Model via Chart-to-Table Pre-training and Multitask Instruction Tuning

### 04 Jan 2024 v1

@@ 1 Introduction
The contributions of this paper can be summarized as follows. 1) We present ChartAssistant, a vision-language model for chart comprehension and reasoning. ChartAssistant is versatile enough to solve various chart-related tasks across a wide range of chart types. 2) We build a chart-specific visual instruction-following benchmark dubbed ChartSFT. ChartSFT surpasses existing chart-based benchmarks with its larger instruction-following data corpus, a broader range of tasks and chart types, and more comprehensive data annotations. 3) Extensive experimental results on various downstream tasks demonstrate that ChartAssistant surpasses the previous SoTA method UniChart [30] by 1.0%, 33.8%, 210.0%, 479.0%, 10.1% performance gain on chart-to-table translation, chart open-ended QA, numerical QA, referring QA and summarization, respectively. Notably, ChartAssistant even outperforms advanced multimodal models such as OpenAI’s GPT-4V(ision) [3] and Google’s Bard [6] on chart data in real scenarios.
@@ 2.2 Chart-specific Vision-Language Model
Some methods modify vision-language models for chart-related tasks. Matcha [25] extends Pix2Struct [19] by integrating mathematical reasoning and chart data extraction tasks, excelling at chart question answering and chart summarization. Unichart [30] undergoes multitask instruction tuning for more chart-related tasks, establishing itself as the most versatile and effective chart vision-language model currently available. However, these methods have limitations. They require task-specific fine-tuning for promising results, risking catastrophic forgetting [44, 5] and sacrificed task generalization. Furthermore, these models struggle with mathematical computations, limiting their effectiveness and range of applicable chart types.
@@ 5.1 Baselines and Evaluation
Baselines. We choose SPHINX [23], ChartLLaMa [7], Unichart [30], Matcha [25], Pix2Struct [19], T5 [35] and Chart-T5 [49] as baselines. ChartLLama and Unichart are the current state-of-the-art models that handles the maximum number of chart tasks and delivers the best overall performance. Besides, Unichart also considers the open-ended QA task. Matcha outperforms previous models in mathematical calculations. Pix2Struct and Donut stands out as an excellent document understanding model. We fine-tune these document models on the train set of the respective evaluation datasets and present the results. T5 is a text-to-text model and needs OCR-based system to extract the data table from the chart image, Chart-T5 is a model modified from T5 for chart-related tasks. We use the results from Unichart [30] for them. Note that ChartAssistant does not have undergone task-specific fine-tuning.
@@
Donut 78.1 29.8 8.6 38.2 87.4 13.1 36.3 6.2
SPHINX 11.3 21.7 6.7 16.6 9.4 9.3 4.4 7.2
Pix2Struct 56.0 30.5 10.3 38.2 85.9 12.7 35.6 5.8
@@ 6 Ablation Study
Ours-D w/o align 87.5 40.4 15.8 35.4 87.9 14.8 52.4 60.3
Ours-D w/o refer 89.2 41.4 16.3 38.6 90.7 16.5 62.2 -
Ours-D 91.3 45.3 16.2 40.2 92.0 17.2 72.1 64.2
@@ A.2 Details of Referring QA in ChartSFT
QA Generation. Following the pipeline used in generating numerical QA templates, we extend its application to the referring QA task. As outlined in table LABEL:tab:referQA_template, we define a total of 114 templates, encompassing questions related to label recognition and mathematical calculations. Note that the x_tick of line and area charts is continuous, therefore, we tailor these templates to accommodate such scenarios.
@@ B.1 Experimental Setups
We begin by conducting alignment pre-training, utilizing the chart-to-table translation task for 65k steps. Following that, we engage in multitask instruction tuning. We employ the Adam optimizer [18] with a scheduled learning rate, where the initial rate is set to 5e-5 for ChartAst-D and 2e-6 for ChartAst-S. The input resolution is established at 448×448, while the maximum length in the decoder is defined as 1536 for ChartAst-D and 2048 for ChartAst-S. After training for 330k steps, we perform testing on multiple downstream tasks. During inference, each task receives an image and a textual instruction as input, and the model generates a textual answer. All training processes are carried out on 8xA100 80GB GPUs.
@@
stage1 w/o arXiv 89.4 40.8 15.7 39.1 91.0 16.6 64.1 52.3
stage2 w/o arXiv 89.7 40.6 14.2 36.9 91.2 16.0 56.7 47.7
Ours-D 91.3 45.3 16.2 40.2 92.0 17.2 72.1 64.2

### 10 Jan 2024 v2

@@ 1 Introduction
The contributions of this paper can be summarized as follows. 1) We present ChartAssistant, a vision-language model for chart comprehension and reasoning. ChartAssistant is versatile enough to solve various chart-related tasks across a wide range of chart types. 2) We build a chart-specific visual instruction-following benchmark dubbed ChartSFT. ChartSFT surpasses existing chart-based benchmarks with its larger instruction-following data corpus, a broader range of tasks and chart types, and more comprehensive data annotations. 3) Extensive experimental results on various downstream tasks demonstrate that ChartAssistant surpasses the previous SoTA method UniChart [31] by 210.0%, 479.0%, 15.4% performance gain on numerical QA, referring QA and ChartQA, respectively. Notably, ChartAssistant even outperforms advanced multimodal models such as OpenAI’s GPT-4V(ision) [3] and Google’s Bard [6] on chart data in real scenarios.
@@ 2.2 Chart-specific Vision-Language Model
Some methods modify vision-language models for chart-related tasks [7, 26] or develop plugin for LLM to understand the chart [44]. Matcha [25] extends Pix2Struct [19] by integrating mathematical reasoning and chart data extraction tasks, excelling at chart question answering and chart summarization. Unichart [31] undergoes multitask instruction tuning for more chart-related tasks, establishing itself as the most versatile and effective chart vision-language model currently available. However, these methods have limitations. They require task-specific fine-tuning for promising results, risking catastrophic forgetting [46, 5] and sacrificed task generalization. Furthermore, these models struggle with mathematical computations, limiting their effectiveness and range of applicable chart types.
@@ 5.1 Baselines and Evaluation
Baselines. We choose SPHINX [23], ChartLLaMa [7], Unichart [31], Matcha [25], Pix2Struct [19], T5 [36] and Chart-T5 [51] as baselines. ChartLLama and Unichart are the current state-of-the-art models that handles the maximum number of chart tasks and delivers the best overall performance. Besides, Unichart also considers the open-ended QA task. Matcha outperforms previous models in mathematical calculations. Pix2Struct and Donut stands out as an excellent document understanding model. We fine-tune these document models on the train set of the respective evaluation datasets and present the results. T5 is a text-to-text model and needs OCR-based system to extract the data table from the chart image, Chart-T5 is a model modified from T5 for chart-related tasks. We use the results from Unichart [31] for them. Note that ChartAssistant only has undergone task-specific fine-tuning in Chart-to-text [15]. In fact, we can repeat the training set of this data set in the training set to avoid task-specific finetuning, but this will make the generalization ability of the model weaker.
@@
Donut 78.1 29.8 7.2 38.2 87.4 13.1 36.3 6.2
SPHINX 11.3 21.7 5.4 16.6 9.4 9.3 4.4 7.2
Pix2Struct 56.0 30.5 10.3 38.0 85.9 12.7 35.6 5.8
@@ 6 Ablation Study
Ours-D w/o align 87.5 40.4 13.7 35.4 87.9 13.9 52.4 60.3
Ours-D w/o refer 89.2 41.4 14.1 38.6 90.7 14.6 62.2 -
Ours-D 91.3 45.3 14.1 40.2 92.0 14.9 72.1 64.2
@@ A.2 Details of Referring QA in ChartSFT
QA Generation. Following the pipeline used in generating numerical QA templates, we extend its application to the referring QA task. As outlined in fig. 16, we define a total of 114 templates, encompassing questions related to label recognition and mathematical calculations. Note that the x_tick of line and area charts is continuous, therefore, we tailor these templates to accommodate such scenarios.
@@ B.1 Experimental Setups
We begin by conducting alignment pre-training, utilizing the chart-to-table translation task for 65k steps. Following that, we engage in multitask instruction tuning. We employ the Adam optimizer [18] with a scheduled learning rate, where the initial rate is set to 5e-5 for ChartAst-D and 2e-6 for ChartAst-S. The input resolution is established at 448×448, while the maximum length in the decoder is defined as 1536 for ChartAst-D and 2048 for ChartAst-S. After training for four epochs for ChartAst-D and only one epoch for ChartAst-S, we perform testing on multiple downstream tasks. During inference, each task receives an image and a textual instruction as input, and the model generates a textual answer. All training processes are carried out on 16xA100 80GB GPUs. ChartAst-S outperforms ChartAst-D and has stronger robustness. This is partly due to the special high-resolution image handling method employed by ChartAst-S, which retains more detailed chart information. Additionally, ChartAst-S incorporates richer pre-training knowledge and the larger model possesses greater robustness.
@@
stage1 w/o arXiv 89.4 40.8 13.8 39.1 91.0 14.5 64.1 52.3
stage2 w/o arXiv 89.7 40.6 12.7 36.9 91.2 13.2 56.7 47.7
Ours-D 91.3 45.3 14.1 40.2 92.0 14.9 72.1 64.2